# Hyper-STTN: Social Group-aware Spatial-Temporal Transformer Network for Human Trajectory Prediction with Hypergraph Reasoning

### v2

@@ Abstract
Predicting crowded intents and trajectories is crucial in various real-world applications, including social robots and autonomous vehicles. Understanding environmental dynamics is challenging, not only due to the complexities of modeling pair-wise spatial and temporal interactions but also the diverse influence of group-wise interactions. To decode the comprehensive pair-wise and group-wise interactions in crowded scenarios, we introduce Hyper-STTN, a Hypergraph-based Spatial-Temporal Transformer Network for crowd trajectory prediction. In Hyper-STTN, crowded group-wise correlations are constructed using a set of multi-scale hypergraphs with varying group sizes, captured through random-walk probability hypergraph spectral convolution. Additionally, a spatial-temporal transformer is adapted to capture pedestrians’ pair-wise latent interactions in spatial-temporal dimensions. These heterogeneous group-wise and pair-wise are then fused and aligned though a multi-modal transformer network. Hyper-STTN outperforms other state-of-the-art baselines and ablation models on several public pedestrian motion datasets.
@@ I Introduction
Human trajectory prediction is a pivotal area in computer vision and robotics research, focusing on predicting agents’ future movements by analyzing their past behaviors. This is vital for many real-world applications such as smart city systems and social robot navigation [1, 2]. Despite its importance, forecasting human trajectories in social settings is complex due to the unpredictable nature and varied patterns of crowd movements. The challenge is intensified by the need to understand the intricate interactions and relations among different social groups, which involve complex cooperation and competitive dynamics, as shown in Fig. 1.
In general, human trajectory prediction is influenced by three factors: individual intrinsic situations, pair-wise and group-wise social interactions, and instantaneous intentions. However, these factors have often lacked high-order interaction descriptions or the ability to reason with heterogeneous features. Individual intrinsic situations exhibit correlations between an agent’s current state and temporal features, where future speed or position can often be accurately decoded from historical data as sequence reasoning processes, as demonstrated in previous work [3, 4]. Conversely, modern forecasting systems face challenges in predicting subjective intentions accurately based on limited information, using machine learning or neural network technologies [5]. Consequently, predictors need to enhance their abilities to describe and understand social latent interactions for better overall performance. This enhancement requires leveraging both group-wise embeddings and pair-wise features of crowd dynamics across spatial-temporal dimensions.
@@
To address the aforementioned issues, current research focuses on estimating complex social influences among humans to infer the uncertainty of human movements for trajectory prediction. For instance, [4, 6, 7, 8] utilize neural networks or attention mechanisms to describe pair-wise social interactions among agents across spatial and temporal dimensions. However, the absence of a graph structure limits the further development of captured features. On the other hand, [7, 9] introduce a spatio-temporal graph to model human movements, where each spatial edge presents the interaction between two agents’ vertices, and temporal edges denotes individual temporal dependencies as intrinsic situation. More recently, [5] constructs the dynamics of pedestrian groups as a hypergraph to capture high-order group-wise interaction in crowds for trajectory forecasting. In this approach, each hyperedge can link more than two vertices for group-wise state representation.
@@
Figure 1: HHI (human-human interaction) illustration: Group-wise HHI captures latent correlations among high-level perspectives on group behaviors, while pair-wise spatial-temporal HHI represents individual influences.
@@
While the state-of-the-art (SOTA) algorithms mentioned above have demonstrated outstanding performance, the insufficient inference of human-human interaction (HHI) continues to limit potential improvements, especially in complex or highly dynamic scenarios. None of the previously mentioned approaches have fully addressed HHI in both group-wise and pair-wise feature dimensions [4], nor have they effectively aligned heterogeneous multimodal dependencies [7]. For instance, the absence of group-wise interaction reasoning may overlook potential intergroup harmonization or conflict [9]. In team sports scenarios like basketball, different defensive strategies may involve distinct subgroups, such as zone defense typically including one offensive player with multiple defensive players, and man-to-man policies often involving pairs of players. Also, inadequate spatial-temporal feature inference can introduce ambiguities in HHI [5].
@@
To address these challenges, we introduce Hyper-STTN, a hypergraph-based spatial-temporal transformer network designed to handle both pair-wise and group-wise underlying social interactions, as illustrated in Fig. 1, across spatial-temporal dimensions. The core concept behind Hyper-STTN is its utilization of an effective HHI and social interaction inference framework using multiscale hypergraphs. The main contributions of this paper can be summarized as follows: (1). We propose a novel human trajectory prediction framework called Hyper-STTN, which constructs crowded dynamics using a set of multiscale hypergraphs to leverage group-wise and pair-wise social interactions in both spatial and temporal dimensions; (2). Hyper-STTN aligns heterogeneous features and interactions using a multi-modal transformer network to avoid confusion in interpretations; (3). Hyper-STTN demonstrates outstanding performance compared with other existing SOTA approaches on publicly available pedestrian trajectory datasets.
@@ II-A Related Works
Earlier efforts formulated the task of predicting human trajectories based on handcrafted models, such as the social force model [10] and Gaussian process [11], which utilized fixed physical or mathematical rules to adapt to environmental dynamics. The inflexibility of traditional approaches could lead to overlooking underlying social interaction and forecasting errors in more complex conditions. Inspired by the advancements in machine learning, many deep learning-based algorithms have been deployed to better understand social interactions among pedestrians [4, 6]. However, previous learning-based methods either could not fully utilize the captured features directly or could not effectively leverage long-term dependencies. Especially, due to the successes of attention mechanisms and transformers in sequence learning and pair-wise feature representation [12, 7, 13], there have been explicit descriptions of social interactions via both spatial and temporal dimensions using a spatio-temporal graph model of crowd movements. While the description of pair-wise interactions has been well addressed with the development of transformers and attention modules, all the aforementioned methods have neglected to collect group-wise latent influences for individuals.
While [5] has adapted a hypergraph-based network to learn group-wise and pair-wise features with the multiscale property of hypergraph generation, GroupNet still struggles with isolating spatial features along the temporal dimension. Therefore, the challenge of aligning both group-wise and pair-wise social interactions with long-term dependencies remains a key problem in human trajectory prediction. In this work, we not only incorporate spatial and temporal pair-wise interactions into hypergraph-based group-wise interaction representation, but also introduce a cross-modal attention mechanism to leverage the heterogeneous multi-modality dependencies, based on a multi-modal transformer [14].
@@ II-B Social Interaction Reasoning
Conventional sequence learning models, such as convolutional networks [15] and recurrent networks [4, 12], follow a hierarchical structure that processes information with a tedious and deep order. The lack of parallelization in traditional frameworks can potentially lead to the gradual disappearance or explosion of information after many layers of computations, especially when representing long-term dependencies. In contrast, transformer networks directly capture element-to-element correlations in parallel. This structure has led to significant advancements in sequence learning and pair-wise interaction representation, as demonstrated in [12].
More recently, some existing works [7, 8] have employed transformers to model pair-wise interactions across both spatial and temporal dimensions in trajectory prediction tasks.Given that agents’ future trajectories are highly correlated with their self-spatial and temporal features, those methods have achieved improvements over the SOTA performance. However, despite the transformer’s ability to highlight the importance of each pair of elements via the attention mechanism, they tend to overlook group-wise features for crowd correlation representation.
Unlike standard graph structure where each edge connects only two vertices, hypergraph structures provide a natural means to represent group-wise interactions using hyperedges, which can involve multiple vertices in the same hyperedge. Hypergraph structures have found applications in various fields such as biology, physics, recommendation system, trajectory prediction, and more [16, 17]. For instance, in nature, the correlation between predator-competitor pairs is not simply linear; multiple predator-competitor pairs are challenging to approximate. Thus, [16] introduces high-order interactions among various biotic communities with a hypergraph to stabilize natural environments. Similarly, [5] employs a hypergraph to understand social interactions for forecasting human trajectories. However, most existing approaches either use pre-defined topology structures for generation or create hypergraphs based on direct distance without considering the influence of data distribution. In our work, Hyper-STTN not only considers the influence of data distribution and group scale to generate a multiscale hypergraph, but also leverages crowd dependencies in spatial-temporal dynamics into the hypergraph structure.
@@ III-A Preliminary
We adapt a set of multiscale hypergraphs $\mathcal{G}=(\mathcal{V,E,T}\mathbf{,W})$ to construct crowd movement dynamics for trajectory prediction inference. Where $\mathcal{V}=\{{v_{\rm 1}},{v_{\rm 2}},\cdots,{v_{\rm N}}\}$ is the set of vertexes that presents total ${\rm N}$ agents in the scenario, and =$\mathcal{E}=\{{e_{\rm 1}},{e_{\rm 2}},\cdots,{e_{\rm M}}\}$ denotes that a set of ${\rm M}$ hyperedges are constructed in the hypergraphs, particularly each hyperedge can link more than two vertexes to form a group, and $\mathcal{T}=\{{\tau_{\rm 1}},{\tau_{\rm 2}},\cdots,{\tau_{\rm H}}\}$ presents $\rm H$ scales of hypergrpahs, and $\mathbf{W}=diag({{w}_{\rm e_{1}}},{{w}_{\rm e_{2}}},\cdots,{{w}_{\rm e_{M}}})\in\mathbb{R}^{M\times M}$ defines a weighted matrix of hypergraphs. Lastly, the incident matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$ of hypergraph $\mathcal{G}$ is defined as follows: $\mathbf{H}(v,e)=1,~{}\text{if}~{}v\in e;~{}\mathbf{H}(v,e)=0,~{}\text{if}~{}v\notin e.$ The diagonal vertex matrix $\mathbf{M}_{v}\in\mathbb{R}^{N\times N}$ and diagonal edge matrix $\mathbf{M}_{e}\in\mathbb{R}^{M\times M}$ are composed by the degree of vertex $d(v)=\sum_{e\in\mathcal{E}}w_{(e)}\mathbf{H}(v,e)$ and the degree of edge $d(e)=\sum_{v\in\mathcal{V}}\mathbf{H}(v,e)$.
Additionally, the adjacency matrix of hypergraph $\mathcal{G}$ can be represented by $\mathbf{A}\in\mathbb{R}^{N\times N}$, whereas $\mathbf{A}=\mathbf{H}\mathbf{W}\mathbf{H}^{\top}-\mathbf{M}_{\rm v}$. Hyper-STTN not only models group-wise interaction using pedestrians hypergraph construction, but also leverages individual spatial and temporal interaction to align both group-wise and pair-wise features. Let $\mathbf{X}\in\mathbb{R}^{N\times T_{i}\times d}$ and $\mathbf{\hat{X}}\in\mathbb{R}^{N\times T_{o}\times d}$ denote the past and future trajectory data of crowd with time sequence $T_{i}=8$ or $T_{o}=12$ and dimension $d=2$. And the position of $i$th agent in time $t$ can be defined as $\mathbf{x}_{i}^{t}=(x,y)$. Lastly, Hyper-STTN describes the pedestrian trajectory distribution $\mathcal{P}$ as follows:
@@
where the $\{\mathbf{X}_{1},\cdots,\mathbf{X}_{N}\}$ and$\{\mathbf{\hat{X}}_{1},\cdots,\mathbf{\hat{X}}_{N}\}$ are total input sequence and output sequence in the scenario with all agents. And $\mathbf{X}_{i}=\{\mathbf{x}_{i}^{-{T}_{i}+1},\cdots,\mathbf{x}_{i}^{0}\}$, $\mathbf{\hat{X}}_{i}=\{\mathbf{{x}}_{i}^{1},\cdots,\mathbf{{x}}_{i}^{{T}_{o}}\}$ are the $i$th agent’s input data or output data.
@@ III-B Hyper-STTN Architecture
Hyper-STTN captures both group-wise and pair-wise interactions to reason out HHI for human trajectory prediction tasks, as illustrated in Fig. 2. Firstly, the spatial-temporal transformer network is developed to infer spatial and temporal dimensional pair-wise features. Meanwhile, the hypergraph convolution neural network is employed in parallel to present crowd group-level dependencies, based on crowded multiscale hypergraph construction. Subsequently, above heterogeneous group-wise and pair-wise crowded interaction matrices are fused through a multi-modal transformer to align the spatial-temporal embeddings into the hypergraph. Finally, the crowd embedding mixture is decoded from a CVAE [18] decoder to forecast future trajectories.
@@ III-C Spatial-Temporal Transformer
Inspired by [19, 7, 1], we have developed a spatial-temporal transformer network to abstract the correlation of agents’ movement along spatial-temporal dimensions. Particularly, spatial or temporal dependencies can be represented as multiple attention maps, as shown in Fig. 6. In the spatial transformer, the input data $\mathbf{X}$ is first fed into a positional encoding (PE) layer for sequence information encoding based on [12]. Subsequently, the input data proceeds in the proper order through a layer normalization block (LN), a multihead spatial attention layer, and a feed-forward network (FFN), as shown in Fig. 4. Additionally, residual connections are deployed to stabilize the network. The core of the spatial-temporal transformer is the attention mechanism, where the calculation of the multihead spatial or temporal attention layer is defined as follows:
@@
Figure 3: Group-wise HHI Representation: i) We construct group-wise HHI with a set of multiscale hypergraphs, where each agent is queried in the feature space with varying ’k’ in KNN to link multiscale hyperedges. ii) After constructing HHI hypergraphs, group-wise dependencies are captured by point-to-edge and edge-to-point phases with hypergraph spectral convolution operations.
Additionally, due to the limitations of data collection, some timestpes’ individual data maybe disappear in temporal transformer, and some pedestrians are also possible to crop up in spatial transformer. Therefore, we adapt the mask attention mechanism [20] to address the issue of series varying, so that length-varying sequence data can be processed by spatial-temporal transformer.
@@
$\mathcal{M}(\rm n,t)=\left\{\begin{aligned} -\infty\quad\quad&\mathbf{X}(\mathrm{n,t})={none}\\ \omega(n,t)\quad\quad&otherwise\\ \end{aligned}\right.$ (4)
The framework of temporal transformer and spatial transformer are illustrated as Fig. 4. The spatial transformer and temporal transformer capture pair-wise spatial interaction $\mathbf{\hat{Y}}_{\rm{S}}\in\mathbb{R}^{N\times T_{i}\times d}$ along each timestep and individual’s temporal interaction $\mathbf{\hat{Y}}_{\rm{T}}\in\mathbb{R}^{N\times T_{i}\times d}$ along each agent.

### v3

# Hyper-STTN: Hypergraph Augmented Spatial-Temporal Transformer Network for Trajectory Prediction

@@ Abstract
Predicting crowd intentions and trajectories is critical for a range of real-world applications, involving social robotics and autonomous driving. Accurately modeling such behavior remains challenging due to the complexity of pairwise spatial-temporal interactions and the heterogeneous influence of groupwise dynamics. To address these challenges, we propose Hyper-STTN, a Hypergraph-based Spatial-Temporal Transformer Network for crowd trajectory prediction. Hyper-STTN constructs multiscale hypergraphs of varying group sizes to model groupwise correlations, captured through spectral hypergraph convolution based on random-walk probabilities. In parallel, a spatial-temporal transformer is employed to learn pedestrians’ pairwise latent interactions across spatial and temporal dimensions. These heterogeneous groupwise and pairwise features are subsequently fused and aligned via a multimodal transformer. Extensive experiments on public pedestrian motion datasets demonstrate that Hyper-STTN consistently outperforms state-of-the-art baselines and ablation models.
@@ I Introduction
Human trajectory prediction is a pivotal research topic in computer vision and robotics, aiming to anticipate agents’ future movements based on their past behaviors. The forecasting capability is critical for many real-world applications such as smart city systems and social robot navigation [1, 2]. However, accurately forecasting human trajectories in social environments remains highly challenging due to the inherent stochasticity and unpredictability, and the diversity of crowd dynamics. These difficulties are compounded by the need to model intricate interactions among individuals and groups, involving both cooperative and competitive social behaviors, as illustrated in Fig. 1.
Human trajectory forecasting is governed by three interrelated factors: individual intrinsic states, pairwise and groupwise social interactions, and instantaneous intentions. However, lots of existing approaches have struggled in modeling higher-order interactions and in reasoning across heterogeneous feature domains. Specifically, individual intrinsic states encode correlations between an agent’s current conditions and temporal patterns, implying future velocities or locations to be inferred from historical observations through sequence reasoning processes, as demonstrated in prior studies [3, 4]. Moreover, contemporary forecasting frameworks are still unable to accurately infer subjective intentions from sparse or ambiguous observations, even when employing advanced machine learning or neural network architectures [5]. Thus, addressing these challenges necessitates enhanced representations of latent social dynamics, achieved through the joint integration of group-level embeddings and pairwise features across spatio-temporal dimensions.
@@
Recently, recent research has increasingly focused on estimating complex social influences among humans to better capture the uncertainty of human movements. For instance, [4, 6, 7, 8] employ neural networks and attention mechanisms to model pairwise social interactions among agents across spatio-temporal dimensions. Nevertheless, the absence of an explicit graph structure constrains the expressiveness of the extracted features. Alternatively, some works [7, 9] incorporate spatio-temporal graphs, in which spatial edges encode interactions between agents and temporal edges represent individual temporal dependencies reflecting intrinsic states. More recently, [10, 5] extended this paradigm by constructing pedestrian group dynamics as a hypergraph to capture high-order group-level interactions for trajectory forecasting, where each hyperedge connects multiple vertices to represent collective group dynamics.
@@
Figure 1: HHI feature illustration: groupwise HHI captures latent correlations among high-level perspectives on group behaviors, while pairwise spatial-temporal HHI represents individual influences.
@@
Despite state-of-the-art (SOTA) algorithms have achieved remarkable performance, inadequate modeling of human-human interaction (HHI) continues to constrain further advances on forecasting accuracy, particularly in highly dynamic environments. None of the above approaches fully model HHI across both group-level and pairwise feature dimensions [4], nor do they effectively align heterogeneous multimodal dependencies [7]. For instance, the lack of group-level interaction reasoning may obscure intergroup coordination or conflict [9]. In team sports such as basketball, different defensive strategies exemplify this challenge: zone defenses typically involve one offensive player interacting with multiple defenders, whereas man-to-man schemes entail predominantly pairwise matchup. Furthermore, insufficient spatial-temporal feature inference can exacerbate ambiguities in HHI modeling [5].
@@
To address the aforementioned challenges, we propose Hyper-STTN, a hypergraph-based spatial-temporal transformer network explicitly designed to model both pairwise and groupwise social interactions across spatial-temporal dimensions, as illustrated in Fig. 1. The central idea of Hyper-STTN is to provide an effective framework for inferring HHI and social dependencies by leveraging multiscale hypergraphs. The main contributions of this work are as follows: (1). We introduce Hyper-STTN, which constructs crowd dynamics using a set of multiscale hypergraphs to jointly capture groupwise and pairwise social interactions in both spatial and temporal domains; (2) Hyper-STTN incorporates a multimodal transformer module to align heterogeneous features and interactions, thereby mitigating interpretational ambiguities; (3). Extensive experiments on publicly available pedestrian trajectory datasets demonstrate that Hyper-STTN outperforms existing SOTA approaches.
@@ II-A Related Works
Early efforts to human trajectory prediction relied on handcrafted models, such as the social force model [11] and Gaussian processes [12], which leverage fixed physical or mathematical rules to abstract environmental dynamics. However, the inherent inflexibility of these traditional methods leads to overlooking latent social interactions and producing inaccurate forecasts under complex conditions. Motivated by advances in machine learning, numerous deep learning–based algorithms have been developed to further express social interactions among pedestrians [13, 14, 4, 6]. However, previous learning-based methods either fail to adaptively incorporate the extracted features or cannot effectively leverage long-term dependencies. Inspired by the success of attention mechanisms and transformers in sequence learning and pairwise feature representation [15, 7, 16], recent work has explicitly modeled social interactions across both spatial and temporal dimensions through spatio-temporal graph representations of crowd movements. While pairwise interactions have been effectively addressed through the development of transformers, aforementioned approaches still neglect to capture groupwise latent influences on individual trajectory.
Despite prior studies [10, 5] have adapted hypergraph-based networks to learn groupwise features, they still struggles to disentangle spatial interactions along the temporal dimension. Consequently, aligning groupwise and pairwise social interactions with long-term dependencies remains a critical challenge in human trajectory prediction. In this work, we address this gap by not only integrating spatial and temporal social interactions from the mask attention mechanism into the hypergraph-based interaction representation, but also introducing the cross-modal attention mechanism [17] to effectively exploit groupwise and pairwise heterogeneous multimodal dependencies.
@@ II-B Social Interaction Reasoning
Conventional sequence learning models, such as convolutional networks [18, 19] and recurrent networks [4, 16, 9], typically adopt hierarchical architectures that process information sequentially through network depth orderly. The lack of parallelization can potentially cause information to gradually vanish or explode across many layers of computation, particularly when capturing long-term dependencies. Conversely, transformer networks directly model element-to-element correlations in a fully parallel manner, leading to substantial improvements in sequence learning and pairwise interaction representation, as exemplified by [15].
More recently, some existing works [7, 8, 13, 14] have employed transformer architectures to encode pairwise interactions across both spatial and temporal dimensions in trajectory prediction tasks. Given that agents’ future trajectories are strongly correlated with their self-spatial and temporal features, those methods have achieved improvements over the SOTA performance via the better latent HHI feature representation. Nevertheless, although transformers can emphasize the importance of each pair of elements through attention mechanisms, they still tend to overlook groupwise features that capture higher-order crowd correlations.
Whereas standard graphs restrict edges to pairwise connections, hypergraphs provide a natural and expressive framework for modeling groupwise interactions, with hyperedges capable of simultaneously linking multiple vertices. Hypergraph models have been applied across diverse domains including biology, physics, recommender systems, and trajectory prediction [20, 21]. For instance, in ecological systems the correlations among predator-competitor groups are rarely linear. Especially, multiple predator-competitor interactions are difficult to approximate with pairwise edges alone. Accordingly, [20] introduced higher-order interactions among biotic communities via hypergraphs to better capture and stabilize natural ecosystems. In a similar vein, [5] employed hypergraphs to capture groupwise social interactions for forecasting human trajectories. However, most existing approaches either rely on pre-defined topologies or construct hypergraphs solely from direct distances, without accounting for the underlying data distribution. In contrast, Hyper-STTN explicitly incorporates both pedestrian distribution effects and group scale to generate multiscale hypergraphs, while simultaneously embedding crowd dependencies across spatiotemporal dynamics within the hypergraph structure.
@@ III-A Preliminary
We construct a set of multiscale hypergraphs $\mathcal{G}=(\mathcal{V,E,T}\mathbf{,W})$ for the representation of crowd movement dynamics. Where $\mathcal{V}=\{{v_{\rm 1}},{v_{\rm 2}},\cdots,{v_{\rm N}}\}$ denotes the set of vertices presenting the $\rm N$} agents in the scenario, and $\mathcal{E}=\{{e_{\rm 1}},{e_{\rm 2}},\cdots,{e_{\rm M}}\}$ is the set of $\rm M$} hyperedges among such vertices. Each hyperedge can link more than two vertices to form the groupwise interaction. The set $\mathcal{T}=\{{\tau_{\rm 1}},{\tau_{\rm 2}},\cdots,{\tau_{\rm H}}\}$ specifies total $\rm H$ multiscale hypergraphs. The weight matrix of each hyperedge is defined as $\mathbf{W}=diag({{w}_{\rm e_{1}}},{{w}_{\rm e_{2}}},\cdots,{{w}_{\rm e_{M}}})\in\mathbb{R}^{M\times M}$. Moreover, the incident matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$ of hypergraph $\mathcal{G}$ is defined as follows: $\mathbf{H}(v,e)=1,~\text{if}~v\in e;~\mathbf{H}(v,e)=0,~\text{if}~v\notin e.$ The diagonal vertex matrix $\mathbf{M}_{v}\in\mathbb{R}^{N\times N}$ and diagonal edge matrix $\mathbf{M}_{e}\in\mathbb{R}^{M\times M}$ are composed by the degree of vertex $d(v)=\sum_{e\in\mathcal{E}}w_{(e)}\mathbf{H}(v,e)$ and the degree of edge $d(e)=\sum_{v\in\mathcal{V}}\mathbf{H}(v,e)$, separately.
Additionally, the adjacency matrix of hypergraph $\mathcal{G}$ is given by $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $\mathbf{A}=\mathbf{H}\mathbf{W}\mathbf{H}^{\top}-\mathbf{M}_{\rm v}$. Notably, Hyper-STTN not only enables the modeling of high-order HHI features via multiscale hypergraphs, but also represents the pairwise dependencies with respect to the lowest-scale hypergraph $\tau_{1}$ that degenerates into a standard graph in the scale-1. Let $\mathbf{X}\in\mathbb{R}^{N\times T_{i}\times d}$ and $\mathbf{\hat{X}}\in\mathbb{R}^{N\times T_{o}\times d}$ denote, separately, the past and future trajectory data of crowd, where $T_{i}=8$ and $T_{o}=12$ are the input and output sequence lengths with the dimension $d=2$. The position of $i$-th agent at time $t$ is expressed as $\mathbf{x}_{i}^{t}=(x,y)$. Eventually, Hyper-STTN describes the pedestrian trajectory distribution $\mathcal{P}$ as:
@@
[…]
@@ III-B Hyper-STTN Architecture
Hyper-STTN captures both groupwise and pairwise interactions to reason out HHI for trajectory prediction, as illustrated in Fig. 2. Specifically, the spatial-temporal transformer network and the hypergraph convolution neural network are leveraged in parallel to encode pairwise HHI and groupwise HHI features, constructing a set of multiscale crowd hypergraphs. Subsequently, the heterogeneous interaction embeddings are fused via the cross-attention from the multimodal transformer. Eventually, forecasting trajectories are then decoded by a CVAE decoder [22].
@@ III-C Spatial-Temporal Transformer
Drawing inspiration from [23, 7, 1, 13], we designed a spatial-temporal transformer network to abstract the correlation of pairwise agents across spatial-temporal dimensions. Whereas spatial and temporal dependencies are represented by the relative importance in the attention maps (Fig. 6). For the spatial transformer, the input sequence $\mathbf{X}$ is first processed by a positional encoding layer to incorporate sequence information [15]. The pre-processed data then passes sequentially through a layer normalization block (LN), a multihead mask attention layer, and a feed-forward network (FFN) (Fig. 4), with residual connections employed to stabilize training. At the core of the spatial-temporal transformer is the attention mechanism, which the vanilla multihead attention mechanism is defined as follows:
@@
Figure 3: groupwise HHI Representation: i) We construct groupwise HHI with a set of multiscale hypergraphs, where each agent is queried in the feature space with varying ’k’ in KNN to link multiscale hyperedges. ii) After constructing HHI hypergraphs, groupwise dependencies are captured by point-to-edge and edge-to-point phases with hypergraph spectral convolution operations.
Additionally, due to limitations in data collection, certain timesteps of individual trajectories may be unrecorded in the temporal transformer, while some pedestrians may appear intermittently in the spatial transformer. To address this variability, a masked attention mechanism [24] is employed, enabling the spatial–temporal transformer to robustly handle length-varying sequence data.
@@
where attention mask matrix $\mathcal{M}$ is defined to handle the issue of varying length data and to encode relative distance feature $\omega(n,t)=f_{fc}(dis(\text{agent-pair}))$ into the attention as follows:
@@
The framework of temporal transformer and spatial transformer are illustrated as Fig. 4. The spatial transformer captures pairwise spatial interactions $\mathbf{\hat{Y}}_{\rm{S}}\in\mathbb{R}^{N\times T_{i}\times d}$ among all agents at each timestep, while the temporal transformer summaries individual temporal interactions $\mathbf{\hat{Y}}_{\rm{T}}\in\mathbb{R}^{N\times T_{i}\times d}$ for each agent.
@@
where the pre-processing method of dataset with the generation of attention mask can be accessed at the project website.